OpenStudy (anonymous):

1.Bought a netbook for $250, rate of 25% a year, what's its worth in 3 years? 2. f(x)=2x+3; g(x)=x^2-9.. Find g[f(2)] 3. g(x)=5x^3+2x^2+ 3 x-4 Please help!

OpenStudy (anonymous):

help!

OpenStudy (lgbasallote):

what's the problem for #3

OpenStudy (anonymous):

find x

OpenStudy (lgbasallote):

it's impossible...with the given info...

OpenStudy (anonymous):

ok thank you, do yhou know any of the others?

OpenStudy (anonymous):

sorry, you

OpenStudy (anonymous):

1) I=PTR/100 I=250*25/100 *3 =250 *3/4= CAN U DO

OpenStudy (anonymous):

its worth in 3 years=250+I

OpenStudy (anonymous):

CAN U DO @KayJanae

OpenStudy (anonymous):

2)G(F(X))=G(X^2-9) KEEP X^2 -9 IN PLACE OF X IN G(X) FUNCTION =2(X^2-9)+3 =?.....

OpenStudy (anonymous):

LATER PLUG VALUE OF 2 IN PLACE OF X ING(F(X)) FUNCTION